OpenStudy (anonymous):

Last one for the night and my homework is done. Woohoo. :) When P dollars is invested at interest rate i, compounded annually, for t years, the investment grows to A dollars, where A=P(1 + i)^t. Find the interest rate if $2490 grows to $3630 in 2 years. Round your answer to the nearest hundredth of a percent.

OpenStudy (anonymous):

3630/2490 = x^2 +2x+1

OpenStudy (anonymous):

go from there

OpenStudy (anonymous):

\[2490(1+i)^2=3630\] \[(1+i)^2=\frac{3630}{2490}=1.4578\]

OpenStudy (anonymous):

\[1+i=\sqrt{1.4578}\] \[i=\sqrt{1.4578}-1\]

OpenStudy (anonymous):

\[i=.2074\] or 20.74% i would love to get that!

OpenStudy (anonymous):

hope this helped you get through the assignment.

OpenStudy (anonymous):

you are really good. thanks man.

OpenStudy (anonymous):

welcome. oh one more thing. if you are going to solve \[(x+1)^2=somthing\] don't multiply out. just take \[x+1=\pm\sqrt{something}\]

OpenStudy (anonymous):

we didn't need the minus here because we wanted a positive answer

OpenStudy (anonymous):

ok. will remember. my head hurts. haha. thanks

OpenStudy (anonymous):

welcome